use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
5x^2 + 8x + 7 = 0

Answer by venugopalramana(3286)

use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
5x^2 + 8x + 7 = 0
DISCRIMINANT=B^2-4AC=8^2-4*5*7=64-140=-VE
COMPLEX ROOTS ...2 CONJUGATE COMPLEX NUMBERS ..NO REAL ROOTS